1/1

Joseph M. Scriven ,

What a Friend We Have in Jesus

What a friend we have in Jesus, All our sins and griefs to bear! What a privilege to carry Ev’rything to God in prayer! Oh, what peace we often forfeit, Oh, what needless pain we bear, All because we do not carry Ev’rything to God in prayer!

Have we trials and temptations? Is there trouble anywhere? We should never be discouraged, Take it to the Lord in prayer: Can we find a friend so faithful Who will all our sorrows share? Jesus knows our ev’ry weakness, Take it to the Lord in prayer.

Are we weak and heavy laden, Cumbered with a load of care? Precious Savior, still our refuge; Take it to the Lord in prayer: Do thy friends despise, forsake thee? Take it to the Lord in prayer; In His arms He’ll take and shield thee; Thou wilt find a solace there.

Blessed Savior, Thou hast promised Thou wilt all our burdens bear May we ever, Lord, be bringing All to Thee in earnest prayer. Soon in glory bright unclouded There will be no need for prayer Rapture, praise and endless worship Will be our sweet portion there.

Verkinto de tiu ĉi Angla poemo estas Joseph M. Scriven (∗1820 – †1886).

Arg-372-726 (2006-08-25 15:58:11)

Tiun originan tekston de la kanto mi trovis en la interreto en / Diesen Urtext des Liedes fand ich im Internetz unter: http: // www. cyberhymnal. org/ htm/ w/ a/ f/ wafwhij. htm .

http://www.poezio.net/version?poem-id=372&version-id=,726